### v1

# FLAP: Fully-controllable Audio-driven Portrait Video Generation through 3D head conditioned diffusion model

Abstract
Diffusion-based video generation techniques have significantly improved zero-shot talking-head avatar generation, enhancing the naturalness of both head motion and facial expressions. However, existing methods suffer from poor controllability, making them less applicable to real-world scenarios such as filmmaking and live streaming for e-commerce. To address this limitation, we propose FLAP, a novel approach that integrates explicit 3D intermediate parameters (head poses and facial expressions) into the diffusion model for end-to-end generation of realistic portrait videos. The proposed architecture allows the model to generate vivid portrait videos from audio while simultaneously incorporating additional control signals, such as head rotation angles and eye-blinking frequency. Furthermore, the decoupling of head pose and facial expression allows for independent control of each, offering precise manipulation of both the avatar’s pose and facial expressions. We also demonstrate its flexibility in integrating with existing 3D head generation methods, bridging the gap between 3D model-based approaches and end-to-end diffusion techniques. Extensive experiments show that our method outperforms recent audio-driven portrait video models in both naturalness and controllability. The source code and models are available at (URL omitted for double-blind reviewing).
Machine Learning, ICML
Figure 1. FLAP offers full controllability by supporting various types of control signals
1 Introduction
Portrait video generation, or zero-shot talking-head avatars, aims to animate a static portrait image using speech audio. Diffusion-based methods have demonstrated high precision in lip synchronization, realistic facial behavior, and natural head movements through specialized modules and network architectures.
However, in practical applications like live streaming for e-commerce or film production, naturalness alone is insufficient. For example, in filmmaking, an avatar should turn its head toward someone before starting a conversation. Additionally, avatars are often required to perform specific motions, such as nodding or head-shaking, as dictated by the film script.
As alternatives, LivePortrait (Guo et al., 2024) and MegActor (Yang et al., 2024) use video as input to generate head motions and expressions based on reference videos. These methods transform the challenge into an expression mapping task. However, by requiring additional video input to specify target expressions and motions, their usability is limited. Earlier GAN-based methods (Burkov et al., 2020; Zhou et al., 2021) separate faces into identity and non-identity components, leveraging identity consistency and pose variation across frames. However, these methods still produce unnatural videos, with restricted pose variations and severe identity loss, resulting in faces that often appear generic. The closest attempt, VASA-1 (Xu et al., 2024b), uses FaceVid2Vid (Wang et al., 2021) and Mega-Portrait (Drobyshev et al., 2022) as its motion-to-video model, while generating the motion input via a diffusion model. Although this approach enhances Mega-Portrait’s control by synthesizing controllable motion inputs, it still relies on the original Mega-Portrait for head rotation control, leading to the same generic face issue.
In contrast, more recent works use diffusion-based models to achieve superior naturalness, but at the cost of reduced controllability. For example, EchoMimic (Chen et al., 2024b), Ani-portrait (Wei et al., 2024) and Follow-Your-Emoji (Ma et al., 2024b) introduce 2D motion fields based on selected landmarks. These landmark based approaches project facial landmarks into 2D image and combines them with noisy latent variables to guide the diffusion process, resulting in three major drawbacks. 1) Projecting landmarks into 2D image space destroys the 3D information, hindering the model’s ability to learn 3D structure. 2) Simply adding 2D landmark images to noisy latents destroys the translation invariance, requiring additional alignment (Ma et al., 2024b; Wang et al., 2024a) between the input landmarks and the reference face. 3) Facial landmarks are suboptimal as control signals for human faces, as they provide vague semantic information and are entangled with identity. Using this vague and entangled information as a primary control signal significantly limits the potential of diffusion models. Furthermore, most end-to-end diffusion-based models (Xu et al., 2024a; Tian et al., 2024; Jiang et al., 2024; Cui et al., 2024) do not support user-specified control signals, ignoring the importance of controllability.
To address this limitation and enhance model usability across various scenarios, we present a novel approach called FLAP. As shown in Figure 1, FLAP can generate natural head movements from audio or replace these movements with user-defined actions, such as turning, tilting, looking around, or even jerking. FLAP supports six degrees of freedom (6DoF) for control over head movements, demonstrating its nuanced understanding of 3D space. In addition to motion control, FLAP can generate lifelike expressions from three sources: audio, user-specified expression coefficients, and coefficients extracted from images or videos. Each of these sources can be used independently or combined to create new expressions. For example, as shown in when no control signal is provided, FLAP generates a talking-head video with natural expressions and mild head movements from audio alone. When expressions are provided by an image with a user-modified head pose, FLAP generates the corresponding talking-head video.
In summary, our contributions are as follows: 1) We introduce FLAP, a diffusion-based framework that produces high controllability while maintaining naturalness and high precision in lip synchronization. 2) To the best of our knowledge, FLAP is the very first work that utilizes 3D head coefficients as diffusion conditions. This not only enhanced its controllability, but also brings flexibility in integrating with existing 3D head generation methods, bridging the gap between 3D model-based approaches and end-to-end diffusion techniques. 3) We propose a Progressively Focused Training scheme for expression-motion decoupling, in which facial expression conditions and head pose are progressively focused and learned. This ensures independent control of facial expression and head motion.
2 Related Works
2.1 Non-diffusion-based Portrait Animation
For non-diffusion-based models, previous approaches (Zhou et al., 2020; Chen et al., 2019) mainly introduce intermediate representation and leverage Generative Adversarial Networks (GANs) (Goodfellow et al., 2020) to generate final images. These methods first train an audio-to-motion model to predict head motion from audio, and then warp head motion into video. For example, AudioDVP (Wen et al., 2020) use BFM (Paysan et al., 2009) parameters as intermediate representation, and predict mouth related parameters to reenact expression. FaceVid2vid (Wang et al., 2021) introduced 3D implicit keypoints to further enhance the representation efficiency for warping and reenactment. CodeTalker (Xing et al., 2023) replace BFM with FLAME (Li et al., 2017) improving expressiveness of keypoints. As for audio-to-motion models, recent works (Fan et al., 2022; Gong et al., 2023) introduced transformer based models to enhance long sequences generalization ability and cross-modal alignment. GeneFace (Ye et al., 2023) proposed a variational motion generator to predict landmark sequence from audio and designed a 3DMM nerf render to solve mean face problem. More recent works like Liveportrait (Guo et al., 2024) introduced a mixed image-video training strategy, with a upgraded network architecture and a much larger trainning dataset . It achieved effectively balance of the generalization ability, computational efficiency, and controllability, yet requires video input as motion source.
2.2 Diffusion-based Portrait Animation
Diffusion models (DMs) (Ho et al., 2020; Song et al., 2020) achieves superior performance in various generative tasks including image generation (Rombach et al., 2022; Ruiz et al., 2023; Shi et al., 2024)and editing (Brooks et al., 2023; Cao et al., 2023), video generation (He et al., 2022; Ma et al., 2024a; Wang et al., 2024b) and editing (Li et al., 2024; Qi et al., 2023; Zhang et al., 2023c). Extensive studies (Zhang et al., 2023a; Li et al., 2025; Hu et al., 2021; Rombach et al., 2022) have proven the feasibility of controlling image generation process with condition signal such as text, landmarks, segmentations, dense poses and encoded images.
Relevant to our task of portrait animation, most recent approaches(Tian et al., 2024; Chen et al., 2024b; Wang et al., 2024a; Xu et al., 2024a; Ma et al., 2024b) employed a dual U-Net architecture similar to AnimateAnyone (Hu, 2024) and incorporate temporal module (Guo et al., 2023) into generation process. Besides, EMO (Tian et al., 2024) also employed face location mask and head speed as control signals using controlnet-based mechanism and attention-based mechanism respectively. Following such a paradigm, EchoMimic(Chen et al., 2024b), V-Express (Wang et al., 2024a) and Follow-your-emoji (Ma et al., 2024b) integrate facial landmark conditions by adopting ControlNet (Zhang et al., 2023a) mechanism. X-portrait (Xie et al., 2024) proposed a ControlNet-based motion control module with local masking and scaling mechanism to achieve fine-grained motion control. Hallo2 (Cui et al., 2024) and Instruct Avatar (Wang et al., 2024c) integrate text condition as a guidance for diffusion process. VASA-1 proposed a motion space DiT (Peebles & Xie, 2023) with eye gaze direction, head-to-camera distance and emotion offset as condition to generate motion latent. For motion-to-video module, it relied on a retrained FaceVid2Vid (Wang et al., 2021) and MegaPortrait (Drobyshev et al., 2022) to render final video result.
Apart from studies on control signals, another branch of existing studies focus on enhancing the alignment between audio, expression and head motion. For example, Hallo (Xu et al., 2024a) proposed a hierarchical audio-driven visual synthesis module to enhance audio-visual correlation. Loopy (Jiang et al., 2024) transformed audio feature and landmark feature into a same latent space to ensure strong dependency between audio and head motion.
2.3 Disentanglement on Talking Head Representation
Disentangled controlling of lip pose, expressions, and head movement is a longstanding task in talking head generation. The main target of this task is to decouple these features of a portrait video, so we can achieve independent manipulation of each. (Siarohin et al., 2019) utilize explicit 2D keypoints to characterize facial dynamics. But these 2D keypoints carry insufficient head pose information in 3D space and carry identity information. Other explicit approaches like 3D face models such as BFM (Paysan et al., 2009) and FLAME (Li et al., 2017) can fit head pose and expression accurately while remaining editable, but they suffer from issues such as identity loss or lack of effective rendering methods. As solutions for expressiveness, implicit approaches such as PC-AVS (Zhou et al., 2021) employs contrastive learning to isolate the mouth space related to audio. PD-FGC (Wang et al., 2023) extended PC-AVS by progressively disentangle identity, head pose, eye gaze direction and emotion after lip pose isolated by PC-AVS. Following this line, EDTalk (Tan et al., 2025) introduce a lightweight framework with orthogonal bases, and reduced training cost without performance degradation. However, the introduce of implicit features requires a preset discrete style label (Wang et al., 2023; Tan et al., 2025) or additional video as input, which reduces style diversity or cause inconvenience.
3 Preliminaries
3.1 Latent Diffusion Model
Latent diffusion models (LDM) (Rombach et al., 2022) are a class of diffusion models that operate diffusion process in latent space, achieving stable and fast generation. At each diffusion time step $t$, the noisy latent $z_{t}$ is fed into the network and performs cross-attention with condition $c$. Then network predicts $\epsilon$ conditioned on the given text or other prompts $c$. The overall training objective can be formulated as follows:
$L_{LDM}=\mathbb{E}_{z_{t},c,\epsilon\sim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,C)\|_{2}^{2}\right]$ (1)
where $\epsilon_{\theta}$ denotes denoising net. In Stable Diffusion (SD), $C$ represents text prompts $c_{text}$, image prompts $c_{img}$ or both $C=[c_{text},c_{img}]$. In the field of portrait animation, condition $C$ includes reference image, audio, context frames, landmarks and other additional information.
3.2 3D Morphable Face Model
The research in 3D morphable face modeling aims to accurately capture, parameterize, and reconstruct human face using statistical models. As a fundamental 3D morphable face model, FLAME (Li et al., 2017) is widely adopted in numerous research works (Daněček et al., 2022, 2023; Qian et al., 2024). FLAME use standard vertex based linear blend skinning (LBS) with blendshapes, with $N_{v}$= 5023 vertices, $k$ = 4 joints (neck, jaw, and eyeballs), and $N_{f}$ = 9976 facets. Formally, FLAME is described by a function
$M(\boldsymbol{\beta},\boldsymbol{\theta},\boldsymbol{\psi})\to(\mathbf{V},\mathbf{F}),$ (2)
which parameterize 3D human face into identity shape $\beta\in\mathbb{R}^{|\beta|}$, facial expression $\psi\in\mathbb{R}^{|\psi|}$, and pose parameters $\theta\in\mathbb{R}^{3k+3}$ for rotations around $k$ = 4 joins (neck, jaw, and eyeballs) and global rotations. Each rotation vector of pose parameters is represented by axis angles ($\in\mathbb{R}^{3}$). Given these parameters, FLAME outputs a 3D mesh with vertices $\mathbf{V}\in\mathbb{R}^{N_{v}\times 3}$ and facets $\in$ $\mathbf{F}\in\mathbb{R}^{N_{f}\times 3}$.
Figure 2: Model architecture of FLAP. The main diffusion net accept a generated 3D parameter input, which can also be modified and biased by user input. For audio only scenarios, FLAP utilizes an audio-to-FLAME module, which generate FLAME coefficients from audio. Our main U-net block consists of 4 layers, namely, head motion layer, spatial attention layer, expression layer and temporal layer. These layers are trained in different stages with different input conditions as we proposed Progressively Focused Training scheme, which will be discussed in Section 5.3
4 Methodology
4.1 Overall Framework
FLAP is build upon Stable Diffusion (SD) and AnimateDiff (Guo et al., 2023) utilizing their weights for initialization. Following recent diffusion-based portrait animation methods (Hu, 2024), we employed a dual U-net architecture which consists of two U-Nets initialized with SD 1.5. In this architecture, one of the U-nets, namely ReferenceNet, works as reference image encoder and, therefore, does not participate in the denoising process. The other U-net, namely DenoisingNet, is responsible for denoising process and generating videos guided by conditions provided by ReferenceNet and other modules.
As shown in Figure 2 , the input of our method consists of three parts: 1) a single portrait image of an arbitrary identity, 2) a speech audio clip from an arbitrary person, and 3) optional user specified control signals. As for user control, our method support various forms of signals. For head motion control, users are allowed to use precise rotation angles around X, Y, and Z axes for each frame to generate desired head movement. In addition, users can provide a video or an image to represent the desired head movements. They can even modify input video motions by changing its rotation angles and generate final movement. With those input, users can freely overlay, modify, and then generate speaker videos with desired head movements. Notably, by combining mild head movements generated from audio with user-controlled movements, our method can finally generate user-specified head actions while still preserving the subtle nodding when speaking, thus maintaining naturalness. If a fix head is needed, user can set head angle to a fixed value to generate talking head videos with fixed pose.
Intuitively, head movements and facial expressions are closely tied with the tone and cadences in speech. To ensure such an interconnection, we bind head movements, facial expressions and lip movement to audio utilizing an audio-to-FLAME module. This module operate to generate natural head movement, expressions and lip movements from audio solely, avoided expression disruption from reference image. The detail of this module will be discussed in Section 4.4. And the visualization of this expression disruption in existing methods will be demonstrate in Section 5.4.
4.2 Controlling Diffusion Model with 3D Head Coefficient as Condition
In recent diffusion-based works like Follow-your-Emoji (Ma et al., 2024b), V-Express (Wang et al., 2024a) and EchoMimic (Chen et al., 2024b), there has been a tendency to project landmark information onto image space, and further serve as the condition to guide diffusion process. Although projected landmark images can be directly applied to control net structure with out extra effort, it leads to a loss of head motions in 3D space. In addition, 2D landmarks carry identity information which need careful design of alignment method across different input images. Problems are likely to arise when there is a large appearance variation in source image and driving video from which landmarks are extracted.
To address above problems, we introduce 3D head coefficients which include head pose coefficients and facial expression coefficients to control diffusion model. In FLAP, our 3D head coefficients are based on FLAME, a parametric 3D head model. We use global rotation $\theta_{globalR}\in\mathbb{R}^{3}$ in Equation 2 as an additional diffusion condition for head control. $\theta_{globalR}$ is a sub-vector of $\theta$, which denotes global head rotation in 6 DoF and is represented by 3 angles of rotation around $x$, $y$ and $z$ axis. For facial expression coefficients, we use $\theta_{jaw}\in\mathbb{R}^{3}$, $\theta_{eyes}\in\mathbb{R}^{6}$, $\psi_{eyelids}\in\mathbb{R}^{2}$ and $\psi_{exp}\in\mathbb{R}^{100}$ in Equation 2 for jaw movement, eye features, eyelids features and expressions. Thus, the proposed 3D head conditions for diffusion process in FLAP can be denotes as
$c_{3Dhead}=[\theta_{globalR},\theta_{eyes},\theta_{jaw},\psi_{eyelids},\psi_{exp}]$ (3)
Overall conditions for diffusion process can be summarize as $C=[C_{3Dhead},c_{ref}]$, where $C_{3Dhead}$ is a sequence of $c_{3Dhead}$ that contains 3D head information varying over time, with each $c_{3Dhead}$ corresponding to a frame in the video. We express this formally as $C_{3Dhead}=\{c_{3Dhead_{1}},c_{3Dhead_{2}},c_{3Dhead_{3}},...,c_{3Dhead_{n}}\}$. For all coefficients mentioned above, we follow the settings of FLAME to open up opportunities to incorporate FLAP with other models in FLAME’s ecosystem such as EMOCA (Daněček et al., 2022), EMOTE (Daněček et al., 2023), and other works in upstream tasks.
The purpose of our DenoisingNet is to generate a video that both matches the 3D head information in $C_{3Dhead}$ and the appearance information in $c_{ref}$. Thus, during training, we extract $c_{3Dhead}$ for every frame in training video using FLAME fitting method (Zielonka et al., 2022). Then train the model to generate video frames given 3D head condition sequence $C_{3Dhead}$ of target frames along with a reference image.Target video is used as supervision. The training objective can be formulated as follow:
$L=\mathbb{E}_{z_{t},c,\epsilon\sim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,C)\|_{2}^{2}\right]$ (4)
where $\epsilon_{\theta}$ denotes DenoisingNet. During inference, as shown in Figure 2, a given portrait image is feed into ReferenceNet to get $c_{ref}$. And a sequence of 3D head condition $\{c_{3Dhead_{1}},c_{3Dhead_{2}},c_{3Dhead_{3}},...,c_{3Dhead_{n}}\}$ is generated by audio-to-FLAME module and then manipulated by user signals, and then feed into DenoisingNet to output a talking head video precisely controlled by 3D head conditions.
Table 1: Quantitative comparisons with existing works. 1 ★ in controlability means only support audio or video driven. 2 ★ indecates support both ways. 3 ★ means head pose and facial expression can be controlled independently. FLAP has 4 ★ since it not only supports all features above, but also supports angle input for nuance head pose manipulation and control.The best result for each metric is in bold, and the second-best result is underlined.
[…]
MCNet video - - 0.4001 7.4600 10.3600 0.2992
EchoMimic ★ ★ video - - 0.5890 6.7896 9.7304 0.3919
EDTalk ★ ★ ★ video - - 0.4159 8.4260 9.2331 0.3018
X-Portrait video - - 0.6544 5.8181 10.2428 0.4795
LivePortrait video - - 0.5474 7.6228 10.4424 0.3987
FLAP (ours) ★ ★ ★ ★ video - - 0.6107 7.7837 9.8302 0.4644
SadTalker audio 163.31 551.94 0.4798 6.9274 8.1210 0.3137
EDTalk ★ ★ ★ audio 166.11 592.20 0.4434 7.2472 8.1161 0.3052
EchoMimic ★ ★ audio 148.91 500.56 0.6116 7.0341 8.6307 0.4329
AniPortrait audio 137.27 581.01 0.6518 7.6117 8.3738 0.4920
Hallo audio 139.54 457.43 0.610 7.5069 8.1786 0.4417
Hallo2 audio 141.54 483.95 0.5993 7.7184 8.0047 0.4241
FLAP (ours) ★ ★ ★ ★ audio 136.54 399.78 0.6254 7.8617 8.2916 0.4592
4.3 Progressively Focused Training for Expression-Motion Decoupling
Training process in previous methods (Tian et al., 2024; Ma et al., 2024b; Hu, 2024; Xu et al., 2024a; Wang et al., 2024a) is structured into two stages, namely image training stage and video training stage. For image training, two frames are randomly extracted from a video: one frame is used as the reference image, while the other serves as the target. With reference image used as input for ReferenceNet, the whole model is trained to make predictions conditioned on given reference image without audio input. In video training stage, the model is trained to predict a continuous sequence of frames conditioned on both audio and reference image. However, 3D head coefficients delivers global pose information and detailed face information at the same time. Thus, when they are introduced as condition, model struggles to learn a correct patterns from massive volume of information within one stage.
Additionally, another significant challenge is learning a decoupled representation of facial expression and head pose. Although FLAME model itself decoupled facial expression from head pose, an unavoidable nature of monocular face 3D reconstruction is that the head has one side facing the camera while the other side is not. Thus, when using popular fitting methods (Zielonka et al., 2022; Retsinas et al., 2024)to reconstruct 3D head coefficients from monocular videos, it’s inevitable that head pose information will leak from expression coefficients. Given a video with only the left side of a talking face, 3D head fitting methods can not get correct coefficients of the other side of this face, leading to inaccuracies and flaws in the representation of the unseen half of face. Therefore, a pattern emerges that if certain facial coefficients appear anomalous, it indicates that this side of the head is not facing the camera. As a result, expression coefficients contain a learnable pattern of head pose information. This makes it easy to train a biased model. Visual results are provided in Section 5.3.
To address above issues, we proposed a Progressively Focused Training strategy (PFT) for training efficiency and expression-motion decoupling. Guided by a basic concept that models tend to learn global features of an image before focusing on the finer details. We divide 3D head conditions in FLAP into two groups: detailed expression conditions $c_{exp}=[\theta_{eyes},\theta_{jaw},\psi_{eyelids},\psi_{exp}]$, and head motion condition $c_{motion}=[\theta_{globalR}]$. Head motion condition is learned in earlier stage, while expression conditions is learned in later stages. Thus, our progressively focused training strategy separate training process into three stages. The first stage is head motion stage, which we conduct image training with all layers trainable. In this stage, $c=[c_{head},c_{ref}]$ is the only provided condition. The model is forced to focus on head pose changes between reference image and target image. The second stage is expression stage. In this stage, the model focus on learning facial expressions which includes eyes information $\theta_{eyes}$ and $\psi_{eyelids}$, jaw information $\theta_{jaw}$ and global expression information $\psi_{exp}$. The overall conditions in this stage can be formulated as $c=[c_{exp},c_{ref}]$, where $c_{exp}=[\theta_{eyes},\theta_{jaw},\psi_{eyelids},\psi_{exp}]$. To avoid coupling of expression and head motion, attention layers related to head motion are frozen. In the third stage, video training is conducted, we follow the settings in recent works (Tian et al., 2024; Xu et al., 2024a).
4.4 Audio to FLAME Module
Most existing diffusion methods trends to generate expressions and appearance jointly in a DenoisingNet with condition $c=[c_{audio},c_{ref}]$. We argue that in training, the existence of $c_{ref}$ can disrupt the learning of $c_{audio}$. From our observation, reference images carry more learnable information of expressions than audio. Since the speaker’s emotion often show minimal variation in short videos. Therefore, the model will be misled to learn a biased pattern by simply copy the expressions of reference image to output video. For example, given a smiling face image and angry speech, model might generate a half-smiling video, fail to align with the angry audio input. More visual results of this observation can be found in Section 5.4.
To avoid $c_{audio}$ being affected by $c_{ref}$, we ultilize an audio-to-FLAME module which generate FLAME-based expression coefficients and head motion coefficients solely from audio. This helps us bind facial expressions and head motion to audio, eliminating $c_{ref}$’s impact on expressions. Another advantage of this is that user can control pose and expression freely by manipulating FLAME coefficients. Benefiting from this architecture, users can also extract coefficients from other videos to accomplish video driven portrait animation tasks. An interactive model viewer111 https://flame.is.tue.mpg.de/interactivemodelviewer.html provided by the authors of FLAME visualize the mapping of FLAME parameters to the mesh output.
Generating head pose and expression from audio in 3DMM coefficients space is a well studied task (Sun et al., 2024; Peng et al., 2023). Our Audio-to-FLAME module follows the designed of (Wen et al., 2020). We retained the network with FLAME coefficients instead of BFM. As FLAME coefficients is portable, the audio-to-Flame module can be replaced by any FLAME coefficients generating methods such as EmoTalk(Peng et al., 2023),EMOTE (Daněček et al., 2023), DiffPoseTalk(Sun et al., 2024). Audio-to-FLAME module is first solely trained and then FLAP is fine-tuned as a whole.
Figure 3: Qualitative comparisons. FLAP achieves the most accurate pose control and highest visual quality. Our method is the only one that can handle scenarios with wide-range of head rotations, complex headpieces and makeup. Notice that our method only uses a 3-dim head angle vector and 117-dim expression vector per frame from driving video, while other methods use raw video as driving signal.
5 Experiments
5.1 Implement Details
Our model is trained on public datasets, including HDTF, CelebV-HQ and VFHQ. All video frames are resized to 512$\times$512. For face reconstruction and tracking, we re-implemented MICA (Zielonka et al., 2022) for video tracking as our tracker to fit FLAME coefficients. The dimension of 3D head condition $\theta_{globalR}$, $\theta_{eyes}$, $\theta_{jaw}$, $\psi_{eyelids}$, and $\psi_{exp}$ is 3, 12, 3, 2, and 100, respectively. In head motion training stage, we use videos with head motion coefficient variance in top 20% and train for 150,000 steps. For facial expression training and video training, we train our model with all data for 150,000 and 50000 steps.
5.2 Evaluation
For audio-driven portrait video generation we compare our method with Sadtalker (Zhang et al., 2023b), EDtalk(Tan et al., 2025), EchoMimic (Chen et al., 2024b), AniPortrait(Wei et al., 2024), Hallo (Xu et al., 2024a), and Hallo2(Cui et al., 2024). Since our model also support video driven portrait animation, we compare our method with MCNet (Hong & Xu, 2023), EDtalk (Tan et al., 2025), Echomimic(Chen et al., 2024b), Liveportrait (Guo et al., 2024) and X-Portrait (Xie et al., 2024) on cross-id video-driven tasks. We evaluate all these methods on a open dataset named CelebV-Text (Yu et al., 2023) which is not used for training for most of the baselines and our model. We randomly sampled 30 videos from CelebV-text to ensure the reliability of the results.
To evaluate the image and video generation quality, we measure Frechet Inception Distances (FID) (Heusel et al., 2017), Frechet Inception Distances (FVD) (Unterthiner et al., 2018) and HyperIQA (Su et al., 2020). For audio-visual alignment, we measure Synchronization-C (Sync-C), Synchronization-D (Sync-D)(Chung & Zisserman, 2016). For Generic Face Image Quality (GFIQA), we measure Face-IQA following (Chen et al., 2024a). Since we evaluate cross identity reenactment in video driven task, there is no GT video to compute FID and FVD.
Figure 4: Effectiveness of Progressive Focused Training (PFT). Expression idle indicates that we use the expression coefficients from a canonical head who has a neutral face and canonical head pose (visualized at the bottom right of expression-idle columns). By swapping original expression with this idle expression, we can explore whether the output head pose is affected by expression coefficients. And thus explore the head pose information leakage from expression coefficients.
As shown in Table 1, FLAP outperforms other methods on most of the metrics in audio driven task. For video driven task, although EDtalk has the best Sync scores, its poor H-IQA and Face-IQA indicates its shortcoming in visual quality. X-portrait achieves the best visual quality, but has the worst lip synchronization among all methods. Our model successfully balance visual quality and lip synchronization, outperforming recent methods in quantitative comparison. It should be highlighted that our method only use a 120-dim FLAME coefficients vector for driving, while all other methods rely on the entire video as their driving signal. For qualitative comparisons, as shown in Figure 3, our method demonstrates the best perceptual quality, motion precision and identity similarity in common scenarios and scenarios with wide-range of head rotations, complex headpieces and makeup.
In summary, our method achieve better results than existing works while none of them support comparable controllability. Our work is the only one that support video-driven, audio-driven, disentanglement and angular control.
Figure 5: Visual result of our method and Hallo (Xu et al., 2024a) animating a smiling portrait with angry tone.
5.3 Ablation Study
We conduct an ablation study to prove the effectiveness of our proposed Progressively Focused Training (PFT) scheme. As is shown in the Figure 4 (the second and third left column), although FLAP can be trained to fit the source pose and expression well with out PFT scheme, its head pose is heavily relied on signals leaked from expression coefficients and fail to decouple them. Once we set its expression coefficients idle, the generated head pose is decreased in range. By contrast, when PFT is applied, the expression change will not affect head pose and achieve disentanglement. In addition, when the model only completes the head motion training, it can generate head motion according to pose source, but is not capable of modifying the expression of reference image. In the last column, when the model completes expression training stage, it’s capable of generating head pose and expression from different source. This decoupling of head pose and facial expression allows for independent control and shows the effectiveness of proposed PFT.
5.4 Effectiveness of Audio-to-FLAME module
Our Audio-to-FLAME module is capable of binding the facial expression to audio, making it reference image irrelevant. As shown in Figure 5, Hallo struggles to generate appropriate expressions since it is confused by the smiling face in reference image and angry tone in the audio. Therefore, it generates a mixed and unnatural face. On the contrary, our model generate expressions solely relies on audio though our Audio-to-FLAME module, and thus generate accurate and natural expressions.
6 Conclusion
In this paper, we present FLAP, a novel approach to zero-shot talking-head avatar generation that integrates explicit 3D head pose and facial expression coefficients into a diffusion-based framework. By enabling independent control over head movements and facial expressions, FLAP significantly improves both the naturalness and controllability of generated portrait videos. Our model demonstrates a nuanced understanding of 3D space, offering flexibility in controlling head movements and generating lifelike expressions from diverse sources, including audio and user-specified coefficients. Extensive experiments show that FLAP outperforms existing methods in terms of visual quality, flexibility and controllability, providing a promising solution for real-world applications.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Appendix
A.1 More Results for Flexibility Analysis
Figure 1 demonstrates the visual results of integrating style controlled Audio-DVP with FLAP. As shown in Figure 1 With audio-dvp, the style of mouth while speaking is changed, a greater degree of jaw opening is generated. Thus, by combining with various existing methods, we can generate a range of speaking styles that we desire.
Figure 1: Visual result of integrating talking style from (Zhang et al., 2023b) into Audio-DVP(Wen et al., 2020) with FLAP. The first row demonstrate original FLAP, while the second row shows the result of utilizing Audio-DVP + style control as a replacement of Audio-to-FLAME module.
Figure 2: Visual results of FLAP framework cooperating with other methods. The first and second rows are head pose and expression source images respectively. The other rows are id reference and generated results.
A.2 Flexibility Analysis
FLAP exhibits a high level of flexibility as it can be integrated with existing models by using their feature vectors as diffusion condition. As an example, As an example, EDtalk (Tan et al., 2025) proposed a method to extract implicit representations of facial expressions, head pose, lip pose and identity information from a single portrait image. Therefore, we can replace our 3D head head conditions in FLAP with feature vectors provided by EDtalk. Figure 2 demonstrates the visual results of retraining FLAP with expression and head pose conditions from EDtalk. The generated results have the same pose as the pose reference image, and the same expression as the expression reference image. Pose and expression features extracted by EDtalk(Tan et al., 2025) are used as intermediate representations.
Further, the replaceable audio-to-FLAME module offers flexibility. As an example, retraining Audio-DVP (Wen et al., 2020) and add style embedding following (Zhang et al., 2023b) can change the style of talking. Visual results are in the supplementary materials.
Figure 3: More portrait animation results. Reference images are from Loopy(Jiang et al., 2024), EchoMimic (Chen et al., 2024b), Follow-Your-Emoji (Ma et al., 2024b) and EMO (Tian et al., 2024) and HDTF dataset.
A.3 More User Controlled Animation Results
We provide more animation results in Figure 3. In the last row, we animate a portrait with audio signal only and set her head pose fixed. In the rest of the rows, we controlled the characters to turn their heads, nod, look upwards, and tilt their heads. In the meantime, they maintain speaking according to the same audio input of the last row. We highlight that all these head pose controls are done by inputting the head rotation angles, no additional video signal is involved.
Figure 4: Our methods avoids the mean face and mean head problem of vasa (Xu et al., 2024b)
A.4 Comparison with VASA
Since vasa (Xu et al., 2024b) is not a open source project, we can only give comparison based on images in its paper. Here we find vasa official demo image suffers severe mean face problem, losing its hair and face feature of the reference image. Vasa model seems tend to generate a face of a average man and fail to capture the detailed feature of reference image. From our observation, this frequently happened in most FaceVid2Vid(Wang et al., 2021) backboned models. In fact, vasa uses a retrained FaceVid2Vid as its image encoder and decoder, which limit its expressiveness. Figure 4 shows our method can successfully avoid this mean face problem.